# Receptogenesis in a Vascularized Robotic Embodiment

### 10 Mar 2026 v1

@@ 2.3 Robotic Demonstration
To demonstrate the efficacy of ex novo hardware fabrication for robots in environmental settings, we developed a representative moth-inspired platform (Fig. 4a). This demonstration highlights the synergy between environmental stimulus (UV-light) and body adaptation via material-level response (UV receptivity), which is expressed with behavioral outputs such as visual signalling and wing flapping (mediated by ”central nervous system” represented by embedded microprocessors).
UV exposure permanently reduced the material’s permeability and unlocked a latent UV-sensing capability, transforming it from a passive substrate into a functional receptive area through which the robot could transduce environmental cues into behavioral outputs. The robot started in minimal configuration, with embedded contacts representing latent positions of the receptors in PETG with nonexistent sensitivity to UV stimulus. The robot’s operation represents co-expression of central strategic level and local material-level execution. Upon a central decision (suggesting the emergent need of UV receptibility in the future, not focused in this work), the robot casts a cocktail of precursors to the vascular system spanning the entire moth-shaped embodiment, vascular shadow area , internal vascular volume of approximately to prepare the embodiment for UV encounter. Internal circulatory fluid reached the entirety of the vascular structure, as seen in the timestamps of dry structure filled with isopropanol (Fig 4c, Supplementary Movie 2) and when adding blue tracer to the isopropanol (Fig 4d, Supplementary Movie 2). The liquid moves forward first in the leading vein and then to the directly connected distributed permeability area, pump advection assisted by wetting (capillary forces) and diffusion, respectively. Our system is resilient to many issues usually found when working with liquid in confined spaces(e.g. bubbles).
At low UV exposure ( LED at ), the prepared (Py-PETG, not PPy-PETG) embodiment did not sense UV (Supplementary Figure S2). Upon encounter of the prepared embodiment with intensive UV stimulus ( LED at ), the embodiment reacted by receptive area synthesis (Figure 4e, Supplementary Movie 3), which led to UV-responsiveness (sensitive to both test and synthesis LED) that completed the control loop by activating wing flapping or visual signalling without contribution by the central control level (Figure 4f, Supplementary Movie 4).
@@ 2.4 Discussion
Chemistry is an undertapped toolkit in robotics that makes available a broader range of bodily modifications than reconfigurations of preassemblies, providing the desired match to real environments. The tight coupling between a chemical reaction and its environment — bridging molecular dynamics and macroscopic structure — provides the basis for natural intelligence as a kind of open-ended computation[18]. Chemical signals are used in multisensory neuromorphic computing [44], augmenting visual sensory data with olfactory context, similarly to biological organisms. Even when narrowly considering PPy, the case study material in this paper, the synthesis result is based on the precursor choice and environmental (synthesis) conditions, leading to wide range of possible outcomes (e.g., towards bioelectronic interfacing [11], ionic sensitivity [42], corrosion resistance [10], or electromagnetic wave shielding [20]). Even considering the light-patterning technique, external stimulus can modulate photocrosslinking patterns for microphase separation of a copolymer, achieving stiffness patterning in a chemically identical material [30], providing possible application scenarios in our vascularized system.
@@
The realization of ex novo receptogenesis through vascularized material transport introduces a paradigm where physical adaptation is governed by molecular-scale interactions. In this framework, the computational burden for bodily modifications is offloaded to the interaction between environmental stimuli and molecular structures, where the synthesis process itself serves as the execution of a program that combines pre-determined bodily affordances (here, the PPy precursos) with situation-specific spatiotemporal information (here, local UV exposure). The newly synthesised hardware is a physical record of the environmental conditions. In the demonstrated case, the wing structure of the robotic lepidopteran transformed from a translucent, non-conductive state to a black, UV-sensitive ”phenotype”. This allows for the augmentation of existing robotic components through site-specific chemical changes, without digital symbolic representation, effectively bridging the groundedness gap between digital control and material reality.
@@
The coordination of advection-based global transport and infusion-based local ”last-mile delivery” caters to a high surface area and volume. At the subsystem level, the system enables high-density functionalization, suggesting a viable pathway towards organism/life-level intertwined hierarchical functionality that, within the current fabrication-centric paradigm, is challenging to fabricate and nearly impossible to modify or repair post-deployment. The high ratio between the total circulatory surface area and the footprint of individual components allows a centralized fluidic reserve to address and modify a vast number of discrete physical features, analogous to a single-wire architecture robustly managing complex electronic arrays. The vascular architecture retains compatibility with microfluidic precision manipulation and targeted delivery approaches; e.g., controlled wetting in the vasculature could target synthesis templated by a central command unit and executed in a distributed manner. Both levels of control - central and distributed - are needed in a practical system, but currently it is either one or the other, without vertical communication, which handicaps performance.
@@
As shown in this study, vascularized composites devised to approach biological levels of adaptability pose challenges, e.g., with tailoring material compositions for suitable robotic preparedness. This challenge is increasingly supported by advancements in computational chemical programming [21] and machine learning [12], which allow for the design and testing of molecular structures and chemical reactions. By leveraging these computational tools, it is possible to design (multiple simultaneous) cascades for stimuli at the specificity similar to at the level of molecular receptors in biology. Materials capable of robot-wide physical actions establish a robust framework for constitutional evolution and the generation of ad hoc hardware in unstructured environments.
@@ 3 Conclusion
In conclusion, we demonstrated the ex novo generation (Figure 1) of UV-sensing hardware via photopolymerization within a moth-inspired, vascularized robotic embodiment. By employing a hierarchical vascularization strategy (Figure 2) - combining global advection defined by 3D printed structure with local infusion - we successfully distributed precursors (namely, pyrrole) across the body, enabling effective volume activation upon local exposure to environmental stimulus (UV-light). The resulting receptive area (PPy) exhibited a quantifiable impedance shift due to reversible (bi)polaron formation upon UV exposure (Figure 3). The effective creation of a UV receptor was successfully demonstrated with visual signalling (LED lights) and wing flapping (actuation) (Figure 4). Our work, realized through conventional chemistry and 3D printing, offers a new paradigm for situated robotics, particularly where environmental adaptiveness and autonomous functional growth are required.
@@ 4 Methods
Electrical characterization. Biologic BP-300 electrichemical impedance analyzer was used to monitor impedance in PEIS mode, constant , and amplitude (Figure 3g). For Supplementary Figure S2, constant (Py-PETG) and constant (PPy-PETG) was used with amplitude, Savitzky-Golay filter (window size 21 and polynomial order 3) was applied with Python for data processing. Oscilloscope (Rigol DS1054) was used to capture transients in Fig 3f.
@@
The magnitude and sign of the impedance rate were indicated using LED‑based signaling. ’Signal detected’ mode was enacted by four successive 4-Hz LED blinks, corresponding to the impedance rate exceeding a threshold, whereas the color of the active LED - red or yellow - indicated its sign (decrease or increase, respectively). A single simultaneous blink of both LEDs indicated the impedance rate was below the threshold. Count-based glitch filtering was implemented using a five-count circular buffer: five consecutive impedance-decrease readings triggered a ’reaction’ mode that activated the wing-flapping mechanism (thorax receptor only), accompanied by the red LED blinking rapidly ten times.
Flapping wing mechanism. Nitinol wire (nitinolxy transition temperature) was woven in 14 passes, each, between the thorax and the wing sections. The binary logic-level control signal from the embedded receptor was applied to the gate of an n-type MOSFET (PHB66NQ03LT) that connected the Nitinol wire to an external power source, causing Nitinol to recover the programmed straight shape using electrothermal activation. One flap consisted of activation at , followed by for natural convective cooling. A recoil spring and rubber bands aided the return to the starting position.
@@
Large Language Models. Microsoft Copilot (GPT-5.2; Microsoft Corporation; accessed December 2025–January 2026) was employed as an assistive tool for embedded microprocessor programming and Python-based script development, including LED intensity analysis and data visualization. All AI-assisted outputs were critically reviewed, validated, and revised by the authors to ensure scientific accuracy and reproducibility.
@@ Acknowledgements
Conceptualization: K.A.P., E.S., I.M.; Data Curation: K.A.P., L.Z.; Funding Acquisition: T.T, A.A., I.M.; Investigation: K.A.P., L.Z., H.P., A.M.; Methodology: K.A.P, L.Z., H.P., U.J.; Project Administration: I.M.; Supervision: T.T., I.M.; Validation: K.A.P., L.Z., I.M.; Visualization: K.A.P., L.Z., I.M.; Writing: K.A.P, L.Z., I.M.; All authors contributed to results discussions and manuscript editing.

### 19 May 2026 v2

@@ 2.3 Robotic Demonstration
To demonstrate the efficacy of ex novo hardware fabrication for robots in environmental settings, we developed a representative moth-inspired platform (Fig. 4a). This demonstration highlights the synergy between environmental stimulus (UV-light) and body adaptation via material-level response (UV receptivity), which is expressed via behavioral outputs such as visual signalling and wing flapping (mediated by ”central nervous system” represented by embedded microprocessors).
UV exposure permanently reduced the material’s permeability to photons and unlocked a latent UV-sensing capability, transforming it from a passive substrate into a functional receptive area through which the robot could transduce environmental cues into behavioural outputs. The robot started in minimal configuration, with embedded contacts representing latent positions of the receptors in PETG with nonexistent sensitivity to UV stimulus. The embodiment consisted of a thorax (vascular shadow area ) and two wings (each with vascular shadow area , approximately 25 times the size of the thorax, highlighting the scalability of the system) with a total internal vascular volume of approximately . The robot’s operation represents the co-expression of the central strategic level and the local material-level execution. Upon a central decision (suggesting the emergent need for UV receptivity in the future, not addressed in this work), the robot casts a cocktail of precursors into the vascular system spanning the entire moth-shaped embodiment. In line with the open circulatory systems of animals that expose the entire vasculature to the same circulatory fluid with minimal control and exert its effect locally, here, the whole vasculature was primed for UV encounter, but a receptor developed in irradiated areas only. Indeed, internal circulatory fluid (here, isopropanol, to demonstrate the compatibility of the same vascular system with multiple liquid carriers) reached the entirety of the vascular structure, as seen in the timestamps of dry structure filled with solvent (Figure 4c, Supplementary Movie 3), and liquid agitation is shown by adding a blue tracer to the solvent (Figure 4d, Supplementary Movie 3). The circulatory liquid passed from the leading vein directly into the distributed permeability (vascularized) area, assisted by pump advection and wetting (capillary forces). Our vascularized system is resilient to many issues, e.g., bubbles, commonly encountered when working with liquids in confined spaces, and uses just a single hydraulic pump with no additional valves or other flow control systems to serve the entire embodiment.
At low UV exposure ( LED at ), the prepared (Py-PETG, not PPy-PETG) embodiment did not sense UV (Supplementary Figure S3). Upon encounter of the prepared embodiment with intensive UV stimulus ( LED at ), the embodiment reacted by receptive area synthesis (Figure 4e, Supplementary Movie 4), which led to UV-responsiveness (sensitive to both test and synthesis LED) that completed the control loop by activating wing flapping (electrothermal activation of nitinol by binary MOSFET-controlled heating at 6.6 W to cause recovery of a preprogrammed closed-wing shape; the flat-wing configuration was recovered by natural convective cooling and spring-assisted return) or visual signalling (Figure 4f, Supplementary Movie 5).
@@ 2.4 Discussion
Chemistry is an undertapped toolkit in robotics: it opens up a broad range of bodily modifications beyond the combinatory rearrangement of preassembled parts and tightly couples those modifications to the unique conditions under which they occur. This coupling between chemical reaction and environment — bridging molecular dynamics and macroscopic structure — has been proposed as a basis for open-ended computation in natural systems [19], and is already being exploited in multisensory neuromorphic platforms that fuse chemical and visual cues [52]. In the same spirit as the present work, Samal et al. recently demonstrated in vivo polymerization of n-doped conducting polymers, catalyzed by heme and whole blood, for optically driven neuromodulation in live animals [42] - a striking demonstration that functional electronic hardware can be synthesized inside a living body from circulating precursors. Our work brings this principle into the engineered domain: a vascularized robotic embodiment whose circulating fluid delivers precursors to sites where they are converted, on demand, into the hardware the robot needs (not necessarily centrally aware of the need nor informed of component generation).
The generality of the approach is supported by the vascular delivery architecture and the ex novo hardware generation concept, neither of which is specific to sensing or material choice (UV or PPy, respectively). Vascular systems in robotics have previously been demonstrated across a range of distinct functions - energy storage and actuation [3, 50], communication [18], thermal regulation [20], and structural solidification via fluid-phase transitions [37] - each representing a separate instantiation of the same delivery architecture. Ex novo hardware generation describes robotic operational capability, not a specific chemistry; prior examples may include in-situ 3D printing [21, 40]. Our work adds to this repertoire by demonstrating the synthesis of a new responsive structure within the body in response to external UV dose rather than relying on internal models, as in 3D printing.
@@
The realization of ex novo receptogenesis through vascularized material transport introduces a paradigm where physical adaptation is governed by molecular-scale interactions. In this framework, the computational burden for bodily modifications is offloaded to the interaction between environmental stimuli and molecular structures, where the synthesis process itself serves as the execution of a program that combines pre-determined bodily affordances (here, the PPy precursors) with situation-specific spatiotemporal information (here, local UV exposure). The newly synthesised hardware is a physical record of the environmental conditions. In the demonstrated case, the wing structure of the robotic lepidopteran transformed from a translucent, non-conductive state to a black, UV-sensitive ”phenotype”. This allows for the augmentation of existing robotic components through site-specific chemical changes, without digital symbolic representation, effectively bridging the groundedness gap between digital control and material reality.
In ex novo hardware chemical generation of an internal component, each synthesis event produces a component shaped by the specific conditions of that encounter - environmental light gradients, local chemistry, body geometry - and this variation is not an experimental shortcoming but a defining feature of the generative architecture. Our design decisions embraced situation-specific, and thus intrinsically non-repeatable, synthesis mechanisms; e.g., we used the relative rate of change rather than the absolute value of impedimetry data as a sensitive area readout in a receptor, precisely because the absolute characteristics of each generated sensor will differ and evolve in time. In conventional sensing, such time- and sample-dependent sensitivity would be undesired for analytical applications. Unlike conventional UV sensors that provide instantaneous, memoryless intensity values, this sensing area acts as an intrinsic leaky integrator. After pulse dosing, the material exhibits a slow () return towards its initial baseline, in part through the extension of receptive PPy chains. The accumulation of new bonds in molecular structure implies that the material’s structural topology becomes a tangible record of the agent’s physical experience; i.e., it instantiates experiential material memory. This temporal response represents an incremental, state-dependent hardware update that bridges the gap between passive sensing and active neuromorphic material computation, characterized by an intrinsic stochasticity component in the final structure. We note that this variability has been exploited as a feature in related work - the same principle of environment-coupled synthesis yielding unique, non-reproducible material signatures has recently been demonstrated as a cryptographic primitive [51].
@@
The coordination of advection-based global transport and infusion-based local ”last-mile delivery” caters to a high surface area and volume. At the subsystem level, the system enables high-density functionalization, suggesting a viable pathway towards organism/life-level intertwined hierarchical functionality that, within the current fabrication-centric paradigm, is challenging to fabricate and nearly impossible to modify or repair post-deployment. The high ratio between the total circulatory surface area and the footprint of individual components allows a centralized fluidic reserve to address and modify a vast number of discrete physical features at once, analogous to a single-wire architecture robustly managing complex electronic arrays. The vascular architecture retains compatibility with microfluidic precision manipulation and targeted delivery approaches; e.g., controlled wetting in the vasculature could target synthesis templated by a central command unit and executed in a distributed manner. Both levels of control - central and local - are needed in a practical system, but currently it is either one or the other, without vertical communication, which handicaps performance.
@@
As shown in this study, vascularized composites designed to achieve biological levels of adaptability pose challenges, e.g., in tailoring material compositions for suitable robotic preparedness. This challenge is increasingly supported by advancements in computational chemical programming [25] and machine learning [12], which allow for the design and testing of molecular structures and chemical reactions. By leveraging these computational tools, it becomes possible to design (multiple simultaneous) cascades for stimuli at a specificity similar to that of molecular receptors in biology. Materials capable of robot-wide physical actions establish a robust framework for constitutional evolution and the generation of ad hoc hardware in unstructured environments.
@@ 3 Conclusion
In conclusion, we demonstrated the ex novo hardware generation (Figure 1) of UV-sensing hardware via photopolymerization within a moth-inspired, vascularized robotic embodiment. By employing a hierarchical vascularization strategy (Figure 2) - combining global advection defined by 3D printed structure with local infusion - we successfully distributed precursors (namely, pyrrole) across the body, enabling effective volume activation upon local exposure to environmental stimulus (UV-light). The resulting receptive area (PPy) exhibited a quantifiable impedance shift due to reversible (bi)polaron formation upon UV exposure (Figure 3). The effective creation of a UV receptor was successfully demonstrated with visual signalling (LED lights) and wing flapping (actuation) (Figure 4). Our work, realized through conventional chemistry and 3D printing, offers a new paradigm for situated robotics, particularly where environmental adaptiveness and autonomous functional growth are required.
@@ 4 Methods
Electrical characterization. Biologic BP-300 electrichemical impedance analyzer was used to monitor impedance in PEIS mode, constant , and amplitude (Figure 3g). For Figure 3f and 3g, supply current (Rigol DP832), of the UV LED was varied (, , , , , ). Each rectangular current pulse was followed by a rest period. In parallel, Biologic BP-300 was used to monitor impedance response (PEIS mode, 115 Hz. Drift was corrected with a linear baseline. For Supplementary Figure S3, constant (Py-PETG) and constant (PPy-PETG) was used with amplitude, Savitzky-Golay filter (window size 21 and polynomial order 3) was applied with Python for data processing. Oscilloscope (Rigol DS1054) was used to capture transients in Supplementary Figure S2.
@@
The magnitude and sign of the impedance rate were indicated using LED‑based signaling. ’Signal detected’ mode was enacted by four successive 4-Hz LED blinks, corresponding to the impedance rate exceeding a threshold corresponding to approx 1% relative impedance drop in our sample, whereas the color of the active LED - red or yellow - indicated its sign (decrease or increase, respectively). A single simultaneous blink of both LEDs indicated the impedance rate was below the threshold. Count-based glitch filtering was implemented using a five-count circular buffer: five consecutive impedance-decrease readings triggered a ’reaction’ mode that activated the wing-flapping mechanism (thorax receptor only), accompanied by the red LED blinking rapidly ten times.
Flapping wing mechanism. Nitinol wire (nitinolxy transition temperature) was woven in 14 passes, each, between the thorax and the wing sections. The binary logic-level control signal from the embedded receptor was applied to the gate of an n-type MOSFET (PHB66NQ03LT) that connected the Nitinol wire to an external power source (Rigol DP832), causing Nitinol to recover the programmed straight shape using electrothermal activation. One flap consisted of activation at , followed by for natural convective cooling. A recoil spring and rubber bands aided the return to the starting position.
@@
Large Language Models. Microsoft Copilot (GPT-5.2; Microsoft Corporation; accessed December 2025–January 2026) and Anthropic Claude (Sonnet 4.6 and Opus 4.7; accessed April - May 2026) were employed as an assistive tool for embedded microprocessor programming and Python-based script development, including LED intensity analysis and data visualization. All AI-assisted outputs were critically reviewed, validated, and revised by the authors to ensure scientific accuracy and reproducibility.
@@ Acknowledgements
Conceptualization: K.A.P., E.S., I.M.; Data Curation: K.A.P., L.Z.; Funding Acquisition: T.T, A.A., I.M.; Investigation: K.A.P., L.Z., H.P., A.M.; Methodology: K.A.P, L.Z., H.P., E.S., U.J. I.M.; Project Administration: I.M.; Supervision: T.T., E.S., I.M.; Validation: K.A.P., L.Z., I.M.; Visualization: K.A.P., L.Z., H.P, E.S., I.M.; Writing - Original Draft: K.A.P, L.Z., I.M.; Writing - Review and Editing: K.A.P., L.Z., H.P., A.M., U.J., T.T., A.A., E.S., I.M.